Refer to Interactive Solution 17.61 to review a me
Answer #1

Since it's a constructive interference:

difference in path length = 47.5 - 33.1 = 14.4 m

So constructive intereference will occur when

14.4 = m*lambda

f = v/lambda

f = m*v/14.4

the lowest frequency will be when m = 1

f = 1*343/14.4 = 23.819 Hz

Since this frequency lies above 20 Hz, it can be heard.

Lowest frequency = 23.82 Hz